# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v3

@@ Abstract
Machine learning has taken a critical role in seismic interpretation workflows, especially in fault delineation tasks. However, despite the recent proliferation of pretrained models and synthetic datasets, the field still lacks a systematic understanding of the generalizability limits of these models across seismic data representing a variety of geologic, acquisition and processing settings. Distributional shifts between different data sources, limitations in fine-tuning strategies and labeled data accessibility, and inconsistent evaluation protocols all represent major roadblocks in the deployment of reliable and robust models in real-world exploration settings. In this paper, we present the first large-scale benchmarking study explicitly designed to provide answers and guidelines for domain shift strategies in seismic interpretation. Our benchmark encompasses over $200$ combinations of model architectures, datasets and training strategies, across three diverse datasets (synthetic and real data) including FaultSeg3D, CRACKS, and Thebe. We systematically assess pretraining, fine-tuning, and joint training strategies under varying degrees of domain shift. Our analysis highlights that commonly used fine-tuning practices can lead to catastrophic forgetting, especially when source and target datasets are distributionally disjoint, and that larger models such as Segformer tend to be more robust to adaptation than smaller architectures. Interestingly, we also find that common domain adaptation methods outperform fine-tuning when the distributional shift is large, yet underperform when source and target domains are similar. Finally, we complement conventional segmentation metrics with a novel analysis based on fault characteristic descriptors, revealing how models absorb structural biases from different training datasets. Overall, we establish a robust experimental baseline to provide insights into the tradeoffs inherent to current fault delineation workflows, and shed light on directions for developing more generalizable, interpretable and effective machine learning models for seismic interpretation. The insights and analyses reported provide a set of guidelines on the deployment of fault delineation models within seismic interpretation workflows.
@@ III Fault Delineation Datasets
Figure 4: Charaterization of the three considered datasets using fault-oriented metrics.
@@
Given that proprietary seismic data is, by definition, not publicly accessible, our study intentionally focuses on open-source datasets to ensure reproducibility and transparency. This choice aligns with established benchmarking practices in machine learning, where the goal is to provide reproducible and extensible frameworks rather than case-specific proprietary analyses. Despite being open-source, the selected datasets span markedly different geological, statistical, and labeling characteristics, including synthetic versus real data, varying signal-to-noise ratios, and both expert and crowdsourced interpretations. These variations provide meaningful heterogeneity for assessing domain shift and model generalization. To illustrate this diversity, Figure 4 presents a comparative visualization of the three datasets using the geometric and topological fault metrics introduced in Section II-F. These metrics quantitatively capture differences in fault density, continuity, and complexity, reinforcing that the chosen datasets represent heterogeneous fault conditions suitable for benchmarking generalization. Below we describe and compare the three datasets used for our benchmarking study in detail.
@@ Acknowledgment
[Uncaptioned image] Jorge Quesada received his B.E. and M.S. degrees from the Pontifical Catholic University of Peru. He joined the Georgia Institute of Technology as a Machine Learning PhD student in the department of Electrical and Computer Engineering in 2021, where he is now part of the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). He is interested in leveraging machine learning and image processing tools to study the mechanisms underlying computer and biological vision.
[Uncaptioned image] Chen Zhou is a Ph.D. student in the School of Electrical and Computer Engineering at the Georgia Institute of Technology. He is currently a Graduate Research Assistant in the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). He is working in the fields of machine learning,and image and video processing. His research interests include trajectory prediction and learning from label disagreement for the applications of autonomous vehicle, and seismic interpretation.
[Uncaptioned image] Prithwijit Chowdhury received his B.Tech. degree from KIIT University, India, in 2020. He joined the Georgia Institute of Technology as an M.S. student in the School of Electrical and Computer Engineering in 2021 and is currently pursuing his Ph.D. as a researcher in The Center for Energy and Geo Processing (CeGP) and as a member of the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). His research interests lie in digital signal and image processing and machine learning with applications to geophysics. He is an IEEE Student Member and a published author, with several works presented at the IMAGE conference and published in the GEOPHYSICS journal.
[Uncaptioned image] Mohammad Alotaibi is a Ph.D. student in the School of Electrical and Computer Engineering at the Georgia Institute of Technology. He is a Graduate Research Assistant in the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). His research focuses on machine learning and image processing, with particular interest in domain adaptation for seismic and medical imaging applications.
[Uncaptioned image] Ahmad Mustafa received his PhD in Electrical and Computer Engineering at the Georgia Institute of Technology. His research interests span deep learning, computer vision, and signal processing for medical and geophysical applications. He served as a Staff Scientist at Occidental Petroleum, leading the research and deployment of large-scale computational models for susbsurface applications. His research has been published in leading, high impact journals. He was awarded the prestigious Roger P. Webb ECE GRA Excellence Award in recognition of his research contributions. He has been involved in various teaching and instructional activities in both academic and industrial settings. His teaching services at Georgia Tech earned him several accolades, such as the Outstanding Online Head TA Award, the ECE GTA Excellence Award and the ECE CREATION Award, respectively.
[Uncaptioned image] Yusufjon Kumakov is currently serving as an Assistant Lecturer at Tashkent State Technical University. He earned his Master’s degree in Petroleum Engineering from the Polytechnic University of Turin, Italy. His primary research interests lie in the application of machine learning techniques to geophysics, with a focus on seismic imaging and signal processing.
[Uncaptioned image] Mohit Prabhushankar received his Ph.D. degree in electrical engineering from the Georgia Institute of Technology (Georgia Tech), Atlanta, Georgia, 30332, USA, in 2021. He is currently a Postdoctoral Research Fellow in the School of Electrical and Computer Engineering at the Georgia Institute of Technology in the Omni Lab for Intelligent Visual Engineering and Science (OLIVES). He is working in the fields of image processing, machine learning, active learning, healthcare, and robust and explainable AI. He is the recipient of the Best Paper award at ICIP 2019 and Top Viewed Special Session Paper Award at ICIP 2020. He is the recipient of the ECE Outstanding Graduate Teaching Award, the CSIP Research award, and of the Roger P Webb ECE Graduate Research Assistant Excellence award,all in 2022. He has delivered short courses and tutorials at IEEE IV’23, ICIP’23, BigData’23, WACV’24 and AAAI’24.
[Uncaptioned image] Ghassan AlRegib is currently the John and Marilu McCarty Chair Professor in the School of Electrical and Computer Engineeringat the Georgia Institute of Technology. In theOmni Lab for Intelligent Visual Engineering and Science (OLIVES), he and his groupwork on robustand interpretable machine learning algorithms, uncertainty and trust, and human in the loop algorithms. The group has demonstrated their work on a widerange of applications such as Autonomous Systems, Medical Imaging, and Subsurface Imaging. The group isinterested in advancing the fundamentals as well as the deployment of such systems in real-world scenarios. He has been issued several U.S.patents and invention disclosures. He is a Fellow of the IEEE. Prof. AlRegib is active in the IEEE. He served on the editorial board of several transactions and served as the TPC Chair for ICIP 2020, ICIP 2024, and GlobalSIP 2014. He was area editor for the IEEE Signal Processing Magazine. In 2008, he received the ECE Outstanding Junior Faculty Member Award. In 2017, he received the 2017 Denning Faculty Award for Global Engagement.He received the 2024 ECE Distinguished Faculty Achievement Award at Georgia Tech. He and his students received the Best Paper Award in ICIP 2019and the 2023 EURASIP Best Paper Award for Image communication Journal. In addition, one of their papers is the best paper runner-up at BigData 2024. In 2024, he co-founded the AI Makerspace at Georgia Tech, where any student and any community member can access and utilize AI regardless of their background.
[…]

### v4

@@ Abstract
Machine learning has taken a critical role in seismic interpretation workflows, especially in fault delineation tasks. However, despite the recent proliferation of pretrained models and synthetic datasets, the field still lacks a systematic understanding of the generalizability limits of these models across seismic data representing diverse geologic, acquisition, and processing settings. In practice, distributional shifts between surveys, limitations in fine-tuning strategies, and inconsistent evaluation protocols remain major obstacles to deploying reliable models in real exploration settings. In this paper, we present the first large-scale benchmarking study explicitly designed to evaluate domain shift strategies for seismic fault delineation. Our benchmark spans over 200 experimental setups combining eight architectures, three datasets (FaultSeg3D, CRACKS, Thebe), and multiple training regimes (individual training, fine-tuning, and joint training). We evaluate performance using three complementary metrics (Dice, Hausdorff Distance, and Bidirectional Chamfer Distance) and analyze both segmentation accuracy and structural fidelity. Our results show that fine-tuning across dissimilar domains can reduce source-domain Dice by up to 75%, demonstrating severe catastrophic forgetting, whereas larger models such as Segformer tend to be more robust to adaptation than smaller architectures. We also find that domain adaptation methods outperform fine-tuning under large distributional gaps but underperform when domains are closely aligned. Finally, we complement conventional metrics with an analysis of fault-network descriptors (length, curvature, sinuosity, segmentation density), revealing the nuances in the interplay between architectural choices and data properties. Overall, this benchmark provides a reproducible foundation for evaluating transferability in seismic fault delineation and offers actionable insights for effective deployment of fault delineation models within seismic interpretation workflows.
@@ III Fault Delineation Datasets
(a) Fault-oriented metrics.
(b) Spectral descriptors.
Figure 4: Charaterization of the three considered datasets using (a) fault-oriented metrics and (b) spectral descriptors.
@@
Given that proprietary seismic data is, by definition, not publicly accessible, our study intentionally focuses on open-source datasets to ensure reproducibility and transparency. This choice aligns with established benchmarking practices in machine learning, where the goal is to provide reproducible and extensible frameworks rather than case-specific proprietary analyses. Despite being open-source, the selected datasets span markedly different geological, statistical, and labeling characteristics, including synthetic versus real data, varying signal-to-noise ratios, and both expert and crowdsourced interpretations. These variations provide meaningful heterogeneity for assessing domain shift and model generalization. To illustrate this diversity, Figure 4(a) presents a comparative visualization of the three datasets using the geometric and topological fault metrics introduced in Section II-F. These metrics quantitatively capture differences in fault density, continuity, and complexity, reinforcing that the chosen datasets represent heterogeneous fault conditions suitable for benchmarking generalization. We also showcase a Uniform Manifold Approximation and Projection (UMAP) embedding of frequency descriptors for all three dataset in Figure 4(b), which reveals three well-separated clusters, confirming systematic acquisition and processing-driven distributional shifts. Below we describe and compare the three datasets used for our benchmarking study in detail.